OpenStudy (anonymous):

Which expression is equivalent to 2(m − 5) + 6? 2m − 1 2m + 22 2m − 4 2m + 16 please help

OpenStudy (anonymous):

Let's simplify step-by-step. 2(m−5)+6 Distribute: =(2)(m)+(2)(−5)+6 =2m+−10+6 Combine Like Terms: =2m+−10+6 =(2m)+(−10+6) =2m+−4

OpenStudy (anonymous):

Answer: =2m−4

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Your Welcome.